# font=small

### v1

@@ 5 Results
Table 4: Number of trade order fills that were adverse and non-adverse.
Next, in Figure 4, one can see the wealth and inventory path for the same random 120 second run of the strategy. Here, again, the left and right figures are under the benchmark and improved simulation environments, respectively. The wealth path is marked-to-market at each time step as in equation (4) and the inventory path changes whenever a trade order fill is received as in equation (3). Here, the market-makers wealth path in the left Figure finishes positive, and in the right Figure it finishes negative, over the same 120 second path. The inventory paths’ show that the benchmark environment traded a lot less than our improved environment, which is mostly due to the additional adverse fills this trader received. The additional adverse fills also significantly outweighed the reduced number of non-adverse fills. And so, it’s quite apparent that these paths now evolve very differently. Lower non-adverse fill probabilities and adverse fills has significantly effected how these processes evolved.
In order to tell the full story, we show a histogram, in Figure 5, of the terminal cash values from all 330 simulation runs of the strategy over this trading day. Each bin indicates how often the strategy paths attained a certain P&L. Here, again, the left and right Figures show the benchmark and improved simulation environments, respectively. One can see that, overall, the strategy performed reasonably well in the benchmark environment and much worse in our improved environment. We believe the performance metric in the benchmark environment is over-inflated, as the market-maker received a lot more non-adverse fills than likely in reality, as well not receiving any adverse fills, which clearly significantly effects the performance. In the right Figure, we can see that the strategy does not perform as well and regularly suffers large losses, relatively speaking. This confirms to us that simulating the performance of a short-term style trading strategy, in particular one involving posting many LOs, is very misleading if it does not track adverse fills and include more accurate non-adverse fill probabilities. This is just as evident in ES, NQ and ZN, where similar results for these assets can be seen in the appendix.
Figure 4: Wealth and inventory processes for the same random path of the strategy under the benchmark (left) and improved (right) simulation environments, respectively.
@@ 6 Conclusion and Future Recommendations
In this paper we simulated the performance of a SOC optimal MM strategy that both included and excluded adverse fills, as well as adjusting the probability of a non-adverse fill. Many MM style strategies often trade passively by posting many LOs throughout the LOB, thus the results here are particularly relevant to these types of strategies, but also to any trading strategy that posts a large amount LOs. Models in mathematical finance often make simplifying assumptions in order to find robust solutions to a problem and while many don’t often effect the overall performance too much, some certainly can. In our setting, modelling the price process and the arrival of MOs independently can create large inaccuracies, which is clear from our simulation results. Specifically, our finding shows that excluding adverse fills and non-adverse fill probabilities can significantly inflate the performance of a trading strategy. Adverse trade order fills are never completely avoidable for a trader who devises a LO posting strategy and are particularly apparent in short-term strategies. Although we can try to predict short-term price movements using methods like the short-term alpha predictor given in Cartea et al., (2015) and Cartea et al., 2018b , large orders often randomly enter the market taking out price levels on a whim. Thus, if a MM or trader is posted at these price levels they will receive adverse fills as there usually isn’t enough time to cancel the orders, regardless of some of the speed advantageous some HFTs may have.
@@
In regard to future research recommendations, we think the results could be improved if a more dynamic MM model were used, where the MM can update the important parameters over time. It is clear from our analysis and the literature that LOB data has a clustering effect, thus this could enable the MM to update their non-adverse fill probabilities, as well as improving the short-term alpha predictors ability to predict adverse moves. Papers like Arroyo et al., (2024) and Maglaras et al., (2022) are some new initial works that aim to estimate fill probabilities using deep learning approaches, which seem to develop some promising ideas for devising a more dynamic model. One should also assess the performance of any short-term alpha predictor used to predict adverse moves, as this is one of the key components of the model in this paper. In the SOC framework, this simple short-term alpha predictor helped bring tractable results, but it is unlikely to be that simple to make such price predictions in reality. Lastly, one could also try to improve the price process modelling aspect, by using a jump-diffusion or Levy-style model. These price processes have been proven to more accurately mimic LOB dynamics rather than the general diffusion models, although sometimes more difficult to apply directly in a SOC model.

### v2

@@ 6 Simulation Results
Table 4: Number of trade order fills that were adverse and non-adverse over 330 simulations throughout the trading on April 24th, 2025.
It is quite apparent that the timing of the trade order fills has changed and many trade order fills occur at very unfavorable prices. Earlier, in our simple trading example, where we gave results in Table 2, we saw there that a overwhelming majority of the fills for the strategy simulation in CL were adverse, as well as for ES, NQ and ZN. This test was also conducted on the exact same trading day as this simulation. In Table 4, we show the number of each type of fill that occurred within our improved simulation environment. We can see that using a much lower non-adverse fill probability of $\rho=0.2$ created a similar distribution of fills between non-adverse fills and adverse fills as in Table 2, and we will soon see that the overall performance has now dropped significantly. This trend also persists in the strategy simulations using ES, NQ and ZN data, as can be seen in the appendix. Based on our empirical evidence, we believe that these results better depict how the strategy would perform in reality. However, we may not want to be overly pessimistic by thinking the SOC MM strategy should have the same number of adverse fills as the basic posting strategy, because the SOC MM model has a short-term alpha predictor, which could potentially predict some of these adverse fills. However, it is highly unlikely that any short-term alpha predictor could anticipate all future adverse fills. The simple predictor in Equation (2) lacks the sophistication needed for such predictions, especially given its simplicity and the likely alpha decay it has undergone due to its public availability. Recall our earlier note on the alpha decay of a publicly disclosed trading strategy in Section 3.2.
Next, in Figure 5, one can see the wealth and inventory path for the same random 120-second run of the strategy. Here, again, the left and right figures are in the benchmark and improved simulation environments, respectively. The wealth path is marked-to-market at each time step as in Equation (4) and the inventory path changes whenever a trade order fill is received as in Equation (3). Here, the MM’s wealth path in the left figure ends positive and at a higher value than in the right figure, both over the same 120 second price path. The inventory paths show that the benchmark environment traded quite differently from our improved environment, which is largely due to the additional adverse fills this trader received, but also the lower non-adverse fill probability. Thus, it is quite apparent that these paths now evolve very differently. Lower non-adverse fill probabilities and adverse fills have significantly affected how these processes evolved.
Figure 5: Wealth and inventory processes for the same random path of the strategy under the benchmark (left) and improved (right) simulation environments.
@@
In order to tell the full story, we show a histogram, in Figure 6, of the terminal cash values from all 330 simulation runs of the strategy over this trading day. Each bin indicates how often the strategy paths attained a certain P&L. Here, again, the left and right figures show the benchmark and improved simulation environments, respectively. One can see that, overall, the strategy performed reasonably well in the benchmark environment and much worse in our improved environment. We believe the performance metric in the benchmark environment is over-inflated, as the MM received a lot more non-adverse fills than likely in reality, as well not receiving any adverse fills, which clearly significantly affects the performance. This confirms to us that simulating the performance of a short-term style trading strategy, in particular one that involves posting many LOs, is very misleading if it does not track adverse fills and include more accurate non-adverse fill probabilities. This is just as evident in ES, NQ and ZN, where similar results for these assets can be seen in the appendix.
@@ 7 Conclusion and Future Recommendations
In this paper we simulated the performance of an SOC optimal MM strategy that included and excluded adverse fills, as well as adjusting the probability of a non-adverse fill. Many MM style strategies often trade passively by posting many LOs throughout the LOB, thus the results here are particularly relevant to these types of strategies, but also to any trading strategy that posts a large amount of LOs. Models in mathematical finance often make simplifying assumptions in order to find robust solutions to a problem, and while many do not often affect the overall performance too much, some certainly can. In our setting, modeling the price process and the arrival of MOs independently can create large inaccuracies, which is clear from our simulation results. Specifically, our findings show that excluding adverse fills and non-adverse fill probabilities can significantly inflate the performance of a trading strategy. Adverse trade order fills are never completely avoidable for a trader who devises a LO posting strategy and are particularly apparent in short-term strategies. Although we can try to predict short-term price movements using methods like the short-term alpha predictor given in Cartea et al., (2015) and Cartea et al., 2018b , these methods are quite simplistic and large orders often randomly enter the market, taking out price levels on a whim. Thus, if a MM or trader is posted at these price levels they will receive adverse fills as there usually isn’t enough time to cancel the orders, regardless of some of the speed advantageous some HFTs may have.
@@
In terms of future research recommendations, we believe that the results could be improved if a more dynamic MM model was used, where the MM could update the important parameters over time. For instance, the model could simulate MOs entering the market with varying sizes, whereas we currently assume that each order has a size of 1. It is clear from our analysis and the literature that LOB data has a clustering effect, thus this could more accurately simulate adverse fills, as well as enabling the MM to update their non-adverse fill probabilities and perhaps improve the short-term alpha predictors ability to predict adverse moves. Another extension could be to adjust the posting strategy in the SOC problem, as discussed in Section 3.2, by spreading out LO postings more for assets that experience rapid price changes, such as NQ. Papers like Arroyo et al., (2024) and Maglaras et al., (2022) are some new initial studies that aim to estimate fill probabilities using deep learning approaches, which seem to develop some promising ideas for devising a more dynamic model. The performance of any short-term alpha predictor used to predict adverse moves should also be assessed, as this is one of the key components of the model in this paper. In the SOC framework, this simple short-term alpha predictor helped bring tractable results, but it is unlikely that it would be that simple to make such price predictions in reality. Lastly, one could also try to improve the price process modeling aspect, by using a jump-diffusion or Levy-style model. These price processes have been proven to more accurately mimic LOB dynamics rather than the general diffusion models, although they are sometimes more difficult to apply directly in a SOC model.
@@ Appendix A Appendix: ES, NQ and ZN Strategy Simulations
Note that the non-adverse fill probability should be much lower in ZN as queue sizes are much larger relative to CL, ES, and NQ while trade sizes are very similar, hence the different distribution and more inflated performance results.